OpenStudy (anonymous):

Find the solution to the initial value problem \[f '(x)=4(\cos u-\sin2 u); f(\frac{ \pi }{6 })=0\]

OpenStudy (anonymous):

So far I have \[f(x)=\sin 4u+2\cos 2u+C\]but when I try to plug in pi/6 for u I'm not getting the correct answer for C which should be 3 (according to the back of the book)

OpenStudy (anonymous):

-3 that is

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

The factor 4 can be ignored. Put it as first factor in f(x): \(f(x)=4( \sin u +\frac{1}{2}\cos2u)\). Oh, now I can see your error: \(f(x)=4\sin x + 2\cos x +C\). You typed sin4x...

OpenStudy (zehanz):

This function will give you C=-3, if you set \(f(\frac{\pi}{6})=0\).

OpenStudy (anonymous):

oh you're right

OpenStudy (anonymous):

LOL thanks alot

OpenStudy (zehanz):

YW! I make these errors myself all the time!

OpenStudy (anonymous):

Good thing you caught this one. I was really confused